Differential Inclusions: Existence of Solutions

Priscilla Macansantos Univ. of the Philippines Baguio, Philippines

We investigate differential inclusions using an integral more general than the Aumann integral, and show that under suitable growth and semicontinuity conditions, the initial value problem has a nonempty solutions set.